# Compressed Image Generation with Denoising Diffusion Codebook Models

### v1

@@ 4 Denoising Diffusion Codebook Models
While DDPM is equivalent to DDCM with $K=\infty$, the first question we address is whether DDCM maintains the synthesis capabilities of DDPM for relatively small $K$ values. We compare the performance of DDPM with that of DDCM using $K\in\{2,4,8,16,64\}$ for sampling from pre-trained pixel and latent space models. We compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) to evaluate the generation performance. In LABEL:app:more_random_gens we report additional metrics and provide qualitative comparisons. For pixel space generation, we use a class-conditional DDM trained on ImageNet $256\times 256$ (Deng et al., 2009; Dhariwal & Nichol, 2021), and apply classifier guidance (CG) (Dhariwal & Nichol, 2021) with unit scale. We use the 50k validation set of ImageNet as the reference dataset, and sample 10k class labels randomly to generate the images. For latent space, we use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) trained on $768\times 768$ images and apply classifier-free guidance (CFG) with scale 3 (equivalent to $w=2$ in (Ho & Salimans, 2021)). As the reference dataset, we randomly select 10k images from MS-COCO (Lin et al., 2014; Chen et al., 2015) along with one caption per image, and use those captions as prompts for sampling.
@@ 5 Image Compression with DDCM
The bit rate of this approach is determined by the size of the codebooks $K$, and the number of sampling timesteps $T$. Specifically, the bit-stream length is given by $(T-1)\log_{2}(K)$. Therefore, the bit rate can be reduced by simply decreasing the number of codebooks, or by using a smaller number of timesteps at generation, e.g., by skipping every other step, or by using specific timestep intervals (see LABEL:app:range_t). In the approach described so far, the length of the bitstream increases logarithmically with $K$, making it computationally demanding to increase the bit rate. For instance, even for $K=8192$, $T=1000$ and $768\times 768$ images our BPP is approximately 0.022. Thus, to produce higher bit rates, we propose to refine the noise selected at timestep $i$ by employing matching pursuit (MP) (Mallat & Zhang, 1993). Specifically, at each step $i$, we construct the chosen noise as a convex combination of $M$ elements from ${\mathcal{C}}_{i}$, gathered in a greedy fashion to best correlate with the guiding residual ${\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}$ (as in Eq. 7). The resulting convex combination involves $M-1$ quantized scalar coefficients, chosen from a finite set of $C$ values taken from $[0,1]$. Therefore, the resulting length of the bit-stream is given by $\smash{(T-1)(\log_{2}(K)M+C(M-1))}$, such that $M=1$ is similar to our standard compression scheme, and the length of the bit-stream increases linearly with $M$ and $C$. We apply this algorithm when the absolute bits number crosses $(T-1)\cdot\log_{2}(2^{13})$. Further details are available in LABEL:app:matching_pursuit.
@@
We evaluate our compression method on Kodak24 (Franzen, 1999), DIV2K (Agustsson & Timofte, 2017), ImageNet 1K $256\times 256$ (Deng et al., 2009; Pan et al., 2020), and CLIC2020 (Toderici et al., 2020). For all datasets but ImageNet, we center crop and resize all images to $512\times 512$. We compare to numerous competing methods, both non-neural and neural, and both zero-shot, fine-tuning based, and training based approaches. For the ImageNet dataset, we use the unconditional pixel space ImageNet $256\times 256$ model of Dhariwal & Nichol (2021), and compare our results to BPG (Bellard, 2018), HiFiC (Mentzer et al., 2020), IPIC (Xu et al., 2024), and two PSC (Elata et al., 2024) configurations, distortion-oriented (PSC-D) and perception-oriented (PSC-P). For all other datasets, we use SD 2.1 $512\times 512$ (Rombach et al., 2022) and compare to BPG, PSC-D, PSC-P, ILLM (Muckley et al., 2023), PerCo (SD) (Körber et al., 2024; Careil et al., 2023), and twoCRDR (Iwai et al., 2024) configurations, distortion-oriented (CRDR-R) and perception-oriented (CRDR-R). PSC shares the same pre-trained model as ours, while PerCo (SD) requires additional fine-tuning. For our method, we apply SD 2.1 unconditionally, as we saw no improvement by adding prompts (see further details in LABEL:app:text_effect). We assess our method for several options of $T$, $K$, $M$, and $C$ to control the bit rate. See further details in LABEL:app:compression_details. We evaluate distortion with PSNR and LPIPS (Zhang et al., 2018) and perceptual quality with FID (Bińkowski et al., 2018). For ImageNet, FID is computed against the entire 50k $256\times 256$ validation set. For the smaller datasets we follow Mentzer et al. (2020) and compute the FID over extracted image patches. Specifically, for DIV2K and CLIC2020 we extract $128\times 128$ sized patches, and for Kodak we use $64\times 64$.
As shown in Fig. 5, our compression scheme dominates previous methods on the rate-perception-distortion plane (Blau & Michaeli, 2019) for lower bit rates, surpassing both the perceptual quality (FID) and distortion (PSNR and LPIPS) of previous methods. For instance, our FID scores are lower than those of all other methods at around 0.1 BPP, while, for the same BPP, our distortion performance is better than the perceptually-oriented methods (e.g., PerCo, PSC-P, and IPIC). However, our method under-performs at the highest bit rates, especially when using SD. This is expected due to the performance ceiling entailed by the pre-trained encoder-decoder of SD (Körber et al., 2024; Elata et al., 2024). The qualitative comparisons in Fig. 4 further demonstrate our superior perceptual quality, where, even for extreme bit rates, our method maintains similarity to the original images in fine details. See LABEL:app:compression_details for more details and results.
@@ 6 Compressed Conditional Generation
we obtain the following result (see proof in LABEL:appendix:cond_compression):
@@
In other words, Proposition 6.1 implies that for the loss ${\mathcal{L}}_{\text{P}}$, increasing $K$ leads to more accurate sampling from the posterior $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$, though this results in longer bit-streams. Thus, as long as we have access to $\nabla_{{\mathbf{x}}_{i}}\log{p_{i}({\mathbf{y}}|{\mathbf{x}}_{i})}$ (or an approximation of it) ${\mathcal{L}}_{\text{P}}$ may serve as a sensible option for solving a compressed conditional generation task with DDCM. Interestingly, we show in LABEL:appendix:compression_private_case that our compression scheme from Sec. 5 is a special case of the proposed compressed conditional generation method, with ${\mathbf{y}}={\mathbf{x}}_{0}$ and ${\mathcal{L}}={\mathcal{L}}_{\text{P}}$.
@@ 6.1 Compressed Posterior Sampling for Image Restoration
Qualitative and quantitative results are reported in Fig. 6. As expected, due to the rate-perception-distortion tradeoff (Blau & Michaeli, 2019), we observe that compressing the outputs of DPS and DDNM harms either their perceptual quality (FID), or their distortion (PSNR), or both. This is while our method achieves superior perceptual quality compared to both DPS and DDNM, including their compressed versions. While our method achieves slightly worse PSNR, this is expected due to the perception-distortion trade-off (Blau & Michaeli, 2018). See LABEL:appendix:zero-shot for more details.
@@ 6.2 Compressed Real-World Face Image Restoration
where $Q(\cdot)$ can be any NR-IQA measure, even a non-differentiable one. In LABEL:app:dmax-ot we explain our choice to set ${\bm{r}}({\mathbf{y}})$ as an MMSE estimator.
@@
The results for the WIDER-Test dataset are reported in Fig. 7 (see LABEL:app:dmax-ot for the other datasets). Our approach clearly optimizes each quality measure effectively and generalizes well according to the $\text{FD}_{\text{DINOv2}}$ scores. This is also confirmed visually, where all of our solutions produce high-quality images with less artifacts compared to previous methods. While our approach shows slightly worse distortion, this is once again expected due to the perception-distortion tradeoff (Blau & Michaeli, 2018).
@@ 7 Discussion
We introduced DDCM, a novel generative approach for DDMs that produces high-quality image samples jointly with their lossless compressed bit-stream representations. We found that DDCM achieves comparable generative performance to DDPM, even when the codebooks are extremely small. We leveraged DDCM to solve several compressed image generation tasks, including image compression and compressed restoration, where we achieved state-of-the-art results. Besides image restoration, our compressed conditional generation framework can be used for any type of diffusion guidance, e.g., for text-conditional generation. We demonstrate this option in LABEL:appendix:classifier-guidance and LABEL:appendix:classifier-free-guidance, introducing new classifier-based and classier-free guidance methods that do not use ${\mathbf{y}}$ for decompression. Moreover, we present in LABEL:app:editing preliminary results for compressed image editing using DDCM, by decompressing an image using a desired edit text prompt.

### v2

@@ 4 Denoising Diffusion Codebook Models
While DDPM is equivalent to DDCM with $K=\infty$, the first question we address is whether DDCM maintains the synthesis capabilities of DDPM for relatively small $K$ values. We compare the performance of DDPM with that of DDCM using $K\in\{2,4,8,16,64\}$ for sampling from pre-trained pixel and latent space models. We compute the Fréchet Inception Distance (FID) (Heusel et al., 2017) to evaluate the generation performance. In App. A we report additional metrics and provide qualitative comparisons. For pixel space generation, we use a class-conditional DDM trained on ImageNet $256\times 256$ (Deng et al., 2009; Dhariwal & Nichol, 2021), and apply classifier guidance (CG) (Dhariwal & Nichol, 2021) with unit scale. We use the 50k validation set of ImageNet as the reference dataset, and sample 10k class labels randomly to generate the images. For latent space, we use Stable Diffusion (SD) 2.1 (Rombach et al., 2022) trained on $768\times 768$ images and apply classifier-free guidance (CFG) with scale 3 (equivalent to $w=2$ in (Ho & Salimans, 2021)). As the reference dataset, we randomly select 10k images from MS-COCO (Lin et al., 2014; Chen et al., 2015) along with one caption per image, and use those captions as prompts for sampling.
@@ 5 Image Compression with DDCM
The bit rate of this approach is determined by the size of the codebooks $K$, and the number of sampling timesteps $T$. Specifically, the bit-stream length is given by $(T-1)\log_{2}(K)$. Therefore, the bit rate can be reduced by simply decreasing the number of codebooks, or by using a smaller number of timesteps at generation, e.g., by skipping every other step, or by using specific timestep intervals (see Sec. B.3). In the approach described so far, the length of the bitstream increases logarithmically with $K$, making it computationally demanding to increase the bit rate. For instance, even for $K=8192$, $T=1000$ and $768\times 768$ images our BPP is approximately 0.022. Thus, to produce higher bit rates, we propose to refine the noise selected at timestep $i$ by employing matching pursuit (MP) (Mallat & Zhang, 1993). Specifically, at each step $i$, we construct the chosen noise as a convex combination of $M$ elements from ${\mathcal{C}}_{i}$, gathered in a greedy fashion to best correlate with the guiding residual ${\mathbf{x}}_{0}-\hat{{\mathbf{x}}}_{0|i}$ (as in Eq. 7). The resulting convex combination involves $M-1$ quantized scalar coefficients, chosen from a finite set of $C$ values taken from $[0,1]$. Therefore, the resulting length of the bit-stream is given by $\smash{(T-1)(\log_{2}(K)M+C(M-1))}$, such that $M=1$ is similar to our standard compression scheme, and the length of the bit-stream increases linearly with $M$ and $C$. We apply this algorithm when the absolute bits number crosses $(T-1)\cdot\log_{2}(2^{13})$. Further details are available in Sec. B.4.
@@
We evaluate our compression method on Kodak24 (Franzen, 1999), DIV2K (Agustsson & Timofte, 2017), ImageNet 1K $256\times 256$ (Deng et al., 2009; Pan et al., 2020), and CLIC2020 (Toderici et al., 2020). For all datasets but ImageNet, we center crop and resize all images to $512\times 512$. We compare to numerous competing methods, both non-neural and neural, and both zero-shot, fine-tuning based, and training based approaches. For the ImageNet dataset, we use the unconditional pixel space ImageNet $256\times 256$ model of Dhariwal & Nichol (2021), and compare our results to BPG (Bellard, 2018), HiFiC (Mentzer et al., 2020), IPIC (Xu et al., 2024), and two PSC (Elata et al., 2024) configurations, distortion-oriented (PSC-D) and perception-oriented (PSC-P). For all other datasets, we use SD 2.1 $512\times 512$ (Rombach et al., 2022) and compare to BPG, PSC-D, PSC-P, ILLM (Muckley et al., 2023), PerCo (SD) (Körber et al., 2024; Careil et al., 2023), and twoCRDR (Iwai et al., 2024) configurations, distortion-oriented (CRDR-R) and perception-oriented (CRDR-R). PSC shares the same pre-trained model as ours, while PerCo (SD) requires additional fine-tuning. For our method, we apply SD 2.1 unconditionally, as we saw no improvement by adding prompts (see further details in Sec. B.5). We assess our method for several options of $T$, $K$, $M$, and $C$ to control the bit rate. See further details in App. B. We evaluate distortion with PSNR and LPIPS (Zhang et al., 2018) and perceptual quality with FID (Bińkowski et al., 2018). For ImageNet, FID is computed against the entire 50k $256\times 256$ validation set. For the smaller datasets we follow Mentzer et al. (2020) and compute the FID over extracted image patches. Specifically, for DIV2K and CLIC2020 we extract $128\times 128$ sized patches, and for Kodak we use $64\times 64$.
As shown in Fig. 5, our compression scheme dominates previous methods on the rate-perception-distortion plane (Blau & Michaeli, 2019) for lower bit rates, surpassing both the perceptual quality (FID) and distortion (PSNR and LPIPS) of previous methods. For instance, our FID scores are lower than those of all other methods at around 0.1 BPP, while, for the same BPP, our distortion performance is better than the perceptually-oriented methods (e.g., PerCo, PSC-P, and IPIC). However, our method under-performs at the highest bit rates, especially when using SD. This is expected due to the performance ceiling entailed by the pre-trained encoder-decoder of SD (Körber et al., 2024; Elata et al., 2024). The qualitative comparisons in Fig. 4 further demonstrate our superior perceptual quality, where, even for extreme bit rates, our method maintains similarity to the original images in fine details. See App. B for more details and results.
@@ 6 Compressed Conditional Generation
we obtain the following result (see proof in App. C):
@@
In other words, Proposition 6.1 implies that for the loss ${\mathcal{L}}_{\text{P}}$, increasing $K$ leads to more accurate sampling from the posterior $p_{0}({\mathbf{x}}_{0}|{\mathbf{y}})$, though this results in longer bit-streams. Thus, as long as we have access to $\nabla_{{\mathbf{x}}_{i}}\log{p_{i}({\mathbf{y}}|{\mathbf{x}}_{i})}$ (or an approximation of it) ${\mathcal{L}}_{\text{P}}$ may serve as a sensible option for solving a compressed conditional generation task with DDCM. Interestingly, we show in Sec. C.2 that our compression scheme from Sec. 5 is a special case of the proposed compressed conditional generation method, with ${\mathbf{y}}={\mathbf{x}}_{0}$ and ${\mathcal{L}}={\mathcal{L}}_{\text{P}}$.
@@ 6.1 Compressed Posterior Sampling for Image Restoration
Qualitative and quantitative results are reported in Fig. 6. As expected, due to the rate-perception-distortion tradeoff (Blau & Michaeli, 2019), we observe that compressing the outputs of DPS and DDNM harms either their perceptual quality (FID), or their distortion (PSNR), or both. This is while our method achieves superior perceptual quality compared to both DPS and DDNM, including their compressed versions. While our method achieves slightly worse PSNR, this is expected due to the perception-distortion trade-off (Blau & Michaeli, 2018). See Sec. C.3 for more details.
@@ 6.2 Compressed Real-World Face Image Restoration
where $Q(\cdot)$ can be any NR-IQA measure, even a non-differentiable one. In Sec. C.4 we explain our choice to set ${\bm{r}}({\mathbf{y}})$ as an MMSE estimator.
@@
The results for the WIDER-Test dataset are reported in Fig. 7 (see Sec. C.4 for the other datasets). Our approach clearly optimizes each quality measure effectively and generalizes well according to the $\text{FD}_{\text{DINOv2}}$ scores. This is also confirmed visually, where all of our solutions produce high-quality images with less artifacts compared to previous methods. While our approach shows slightly worse distortion, this is once again expected due to the perception-distortion tradeoff (Blau & Michaeli, 2018).
@@ 7 Discussion
We introduced DDCM, a novel generative approach for DDMs that produces high-quality image samples jointly with their lossless compressed bit-stream representations. We found that DDCM achieves comparable generative performance to DDPM, even when the codebooks are extremely small. We leveraged DDCM to solve several compressed image generation tasks, including image compression and compressed restoration, where we achieved state-of-the-art results. Besides image restoration, our compressed conditional generation framework can be used for any type of diffusion guidance, e.g., for text-conditional generation. We demonstrate this option in Secs. C.5 and C.6, introducing new classifier-based and classier-free guidance methods that do not use ${\mathbf{y}}$ for decompression. Moreover, we present in Sec. C.7 preliminary results for compressed image editing using DDCM, by decompressing an image using a desired edit text prompt.